OpenStudy (anonymous):

I am doing some math and have no clue what this is or they answer help please. a-4 over 5 =12 but the a-4 is above th 5 and well yea

OpenStudy (anonymous):

Equation:\[\frac{ a-4 }{ 5 }=12\]Simply multiply both sides by 5 to get rid of the 5 in the denominator:\[a-4 = 60\]Add 4 to both sides to isolate a:\[a=64\]

OpenStudy (anonymous):

could you help me with this question ohnston High School has a total of 82 boys and girls who play sports. If the number of boys is 16 more than twice the number of girls, how many boys play sports at this high school?